# https://www.projectaria.com/datasets/nymeria

### v1

@@ 4.1 Full-body motion capture and retargetting
where $\mathbf{v}^{i}$ is the local offset of the $i$th landmark defined on our model, and $T^{i}$ is the global transformation of its parent joint. We initialize $\mathbf{v}^{i}$ by manually placing them on the model. Section 0.B.2 provides more details about our human model and motion retargetting.
4.2 6DoF localization and mapping
Data recorded at the same location are globally aligned into a single metric 3D world via Project Aria MPS [7], which employs state-of-the-art visual inertial odometry (VIO), SLAM and mapping algorithms [61, 63, 26]. In a nutshell, first SLAM is run for each individual recording independently. Subsequently, the resulting maps are loop-closed and jointly optimized via visual-inertial bundle adjustment. The output are highly accurate 1KHz trajectories as shown in Fig. 5 (cf. more examples with Fig. 13 in Appendix) – allowing for example to visualize head- and wrist-motion-clusters respectively.
@@
Given the device trajectories, we then align the body poses into this same frame of reference by correlating the dead-reckoned trajectory estimated by XSens. Since the latter accumulates significant drift, there exists no static global transformation to align them. Instead, we assume a constant transformation $T_{HD}$ between the user’s head segment from XSens $H$ and the Aria device $D$ (Aria is firmly held in place with straps and participants are asked to avoid adjusting the glasses during the recording). We then cut the trajectory into a large number of short 4.2 ms segments, and solve the following optimization problem
@@
Figure 6: XSens skeleton projected into observer views. We sample poses from a 20min recording spread over 1.5Km moving trajectory, and project the localized XSens skeleton into the observer RGB view for an end-to-end qualitative assessment of multi-device localization, motion registration and time synchronization. Note the subject is wearing XSens suit under a normal coat.
4.3 Language description of in-context motion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Motion narration of a 5-sec segment. “C squats to pick up the badminton bag and the soccer ball. C extends both arms straight towards the soccer ball and the badminton bag as he squats on the floor. C takes a step forward with his right foot, then bends both legs and slightly positioned both feet on tiptoe as he squats on the floor. C focuses on the soccer ball and the badminton bag.” Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Atomic action of two 5-sec segments. “C lifts his right leg to prepare for a kick.” and “C kicks the soccer ball with his left feet”. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Activity summarization of a 30-sec segment. “C bikes on the road with a colleague.”
Figure 7: Example multi-level annotations. From top to bottom, we show results for motion narration, atomic action and activity summarization.
@@ 4.3 Language description of in-context motion
We define three annotation tasks to cover different levels of detail while scaling up human efforts in a meaningful way. The finest level is motion narration, with the task to describe detailed body posture, e.g. motion direction, velocity, how the upper body and the lower body interact with other people and the environment, the focusing attention etc. The next level annotation is atomic action. Compared to motion narration with elaborate descriptions, annotators are encouraged to use action verbs whenever possible, e.g. describing with “dancing” as opposed to “swing both arms while rotating the body to the right with legs slightly apart”. Last, we define activity summarization. As opposed to previous tasks which create snippets of 3-5 seconds, summarization breaks video into clips of 15-30 seconds to write one sentence summary of the high-level activity. Figure 7 provide examples of each annotation task. The figure shows the benefit of having all three synchronized views - where the egocentric view provides closed-up observations of the hand-object interactions and the environment, the third-person and motion rendering help annotators grasp a holistic understanding of the actions.
4.4 Data processing summary
We manually inspect the XSens data to identify recordings with poor quality. We achieve 92% good-quality rate, which composes the 1200 recordings. The failure cases are mainly due to manual operational errors e.g. inaccurate body measurements, calibration and cable connections. We observe that MPS is highly robust with only a small fraction of recordings losing track intermittently, i.e. 0.38% headset and 6.3% wristband recordings. The majority of cases resume tracking shortly. Note that the fast motion and significant occlusion occurring from the wristband viewpoint is expected to cause a higher error rate. The breakdown statistics by scenario and location is presented in Tab. 7 and LABEL:tab:location-breakdown in the appendix. For language descriptions, we annotated 38.6 hours motion narration, 207 hours atomic action and 196 hours activity summarization. The average video segment is 5 seconds for narration and atomic action and 30 seconds for summarization. In total, the dataset provides 310.5K sentences 8.64M words from a vocabulary of 6545 distinctive words. Note the average word per sentence is 27.8, which is significant longer than existing motion-language narrations. Figure 8 visualizes the language distribution.
@@ 5 Benchmark Tasks and Baselines
The large amount of diverse and contextual motion data we collected offer unprecedented opportunities in building solutions to motion tracking, synthesis, and understanding tasks, with unique multimodal egocentric sensory input and rich annotation supervisions. We present baseline results on three common problems, but they only scratch the surface of our dataset’s potential.
5.1 Motion generation from sparse inputs
When a user is wearing a VR headset and optionally holding controllers, or wearing a pair of AR glasses and optionally wearing wristbands, we would like to reconstruct their full body movements from onboard sensors alone. This is referred to as the "one-point" [55, 48] or "three-point" [16, 40, 25] tracking problem. Existing work can only rely on synthetic sensor input derived from motion capture datasets. Nymeria provides the first large-scale dataset of real sensor input and body motion ground truth for both model training and quantitative evaluation. It also provides real IMU data from head and wrist devices, in contrast to datasets like TotalCapture [87] or DivaTrack [90] where only IMU data from the XSens mocap system are available. As a result, future research could augment the 6D device poses with IMU and other onboard sensors [90], and/or utilize the camera views [48, 54], as well as the rich annotations in our dataset as contextual conditions for egocentric motion generation.
As baselines, we evaluate three state-of-the-art models: AvatarPoser [40] (three-point regression model), BoDiffusion [16] (three-point diffusion model), and EgoEgo [48] (one-point diffusion model). Table 4 evaluates the reconstruction error of AvatarPoser with real and synthetic three-point input. We can see a small but noticeable quality degradation when real sensor input is used. Table 4 evaluates the two diffusion-based models additionally for their motion quality, by measuring the motion FID score [33].
MPJPE Hand Vel.
Mean Lower Upper PE Err.
Real 7.970 16.735 3.125 6.252 16.705
Synthetic 7.309 15.968 2.508 3.465 16.633
Table 3: The full-body three-point tracking baseline AvatarPoser [40] using the real versus synthetic input.
MPJPE FID
@@ 5.1 Motion generation from sparse inputs
EgoEgo 13.216 19.032 10.002 5.14
BoDiffusion 7.980 15.268 5.279 2.32
Table 4: Diffusion-based motion generation baselines with one-point EgoEgo and three-point BoDiffusion.
5.2 Generative motion prior
The amount of data and the coverage in our dataset is ideal for building a foundation model of human motion. Prior attempts of learning pose [68, 84] or motion [77, 51] manifolds rely on the AMASS [58] dataset with around 40 hours of isolated motions, and with a heavy focus on locomotion or professional performance. In contrast, Nymeria consists of 300 hours of daily activities from everyday people as well as information of their surroundings when in action. Our motion distribution leads to a more suitable prior model for downstream applications that concern about everyday human activities.
Here, we provide an ablation of training a Vector-Quantized Variational Autoencoder (VQ-VAE) [64, 24] of motion data in Tab. 6 (cf. details in the supplementary). We compare the number of codebooks used for product quantization (PQ), codebook (CB) size, and latent dimension (Dim) against mean per-joint position error (MPJPE), Procrustes-Aligned (PA-)MPJPE, and joint position acceleration (ACC) errors. The trained encoder can be used as a "motion tokenizer" akin to a text tokenizer for language models to compress input motions as discrete codes [38], or to project noisy motion into the manifold of natural motion [77]. The resulting latent code is useful for autoregressive motion prediction in a generative motion model, with or without conditions [53]; or in a language model with text tokens for motion understanding tasks [38, 32]. When trained well, the decoder can successfully recover full motion with minimal signal loss.
@@ 5.2 Generative motion prior
Table 5: Ablation of motion VQ-VAE [$mm$]. Product quantization contributes the most to performance improvement, followed by increasing the codebook size while decreasing the latent dimensions.
@@
Table 6: Quantitative results of motion understanding. Lacking natural language prior, TM2T performs worse than MotionGPT, which adopts T5 [75] as its language backbone. It indicates the complexity and diversity of our narration.
5.3 Motion understanding in natural language
An especially valuable type of annotation in our dataset is the high quality hierarchical narrations. We provide 38.6 hours of dense motion narrations, 207 hours of atomic action and 196 hours of activity summarization in English. Compared with HumanML3D [31], our narrations are in longer natural sentences, and come with contextual descriptions of objects and environments. It can support both texture to motion generation and motion to text description tasks, at different levels of detail. More importantly, the contextual descriptions are not only paired with the human motion, but also with videos, point clouds, and other sensory data and annotations. The corroboration of both 2D and 3D environment information with language offers exciting opportunities in grounding language and motion research in the physical world.
As baselines, we present a collection of metrics on training and testing MotionGPT [38] and TM2T [32] for the motion to text task in Tab. 6. Specifically, we report BERT score [97], BLEU score [67], CIDEr score [88] and ROUGE-L score [49]. Please refer to the supplementary for the model and training details.
6 Discussions and Conclusions
@@
We propose the Nymeria dataset of unprecedented scale to accelerate the research of egocentric motion understanding. The dataset is the world’s largest collection of human motion in the wild with 300 hours daily activity, 260M body poses of 264 participants across 50 locations. We provide accurate 6DoF tracking, 3D scene representation and gaze estimation where all modalities are synchronized and aligned into a single metric coordinates. Collectively, the dataset captured 399Km of travel by the participants for a total of 201.2M egocentric images, 11.7B IMU samples and 10.8M gaze point. The Nymeria dataset also stands out as the world largest motion-language dataset with 310.5K sentences in 8.64M words text descriptions distributed in 38.6 hours of fine-grained motion narration, 207 hours atomic actions and 196 hours activity summarization.
Limitations.
The mocap suit and miniAria wristbands lead to unnatural visual appearance which might limit the usefulness of videos for computer vision research. The setup also leads to some restriction which directly effect the range of motion. The XSens data quality is known to be affected by motion calibration and body measurements, the accuracy of which can be inconsistent across users. We are aware that our dataset only covers a portion of common daily activities. Important missing activities are the ones in public location, e.g. shopping and highly dynamic motions e.g. driving and fast-moving sports for which our solution has limitations.
Social impact.
@@
Understanding egocentric full-body motion is crucial towards building contextualized AI. However the dataset and algorithms heavily involve personal data. While constructing a valuable dataset to push the boundary of technology, we make best effort respect privacy via data consent, de-identification and minimum data retention.

### v2

@@ 3.2 Data processing
where $\mathbf{v}^{i}$ is the local offset of the $i$th landmark defined on our model, and $T^{i}$ is the global transformation of its parent joint. We initialize $\mathbf{v}^{i}$ by manually placing them on the model. The supplementary provides more details about our human model and motion retargetting.
@@
3.2.2 6DoF localization and mapping with global alignment.
Data recorded at the same location are globally aligned into a single metric 3D world via Project Aria MPS [8], which employs state-of-the-art visual inertial odometry (VIO), SLAM and mapping algorithms [74, 76, 29]. In a nutshell, first SLAM is run for each individual recording independently. Subsequently, the resulting maps are loop-closed and jointly optimized via visual-inertial bundle adjustment. The output are highly accurate 1KHz trajectories (cf. Fig. 4 and supplementary) – allowing for example to visualize head- and wrist-motion-clusters respectively.
Given the device trajectories, we align the body poses into this same reference coordinates by correlating the dead-reckoned trajectory from XSens. Since the latter accumulates significant drift, there exists no static global transformation to align them. Instead, we assume a constant transformation $T_{HD}$ between the user’s head segment from XSens $H$ and the Aria device $D$ (Aria is firmly held in place with straps and participants are asked to avoid adjusting the glasses during the recording). We then cut the trajectory into a large number of short 4.2 ms segments, and solve the following optimization problem
@@
Figure 5: End-to-end quality assessment. We uniformly sample a 20min recording over 1.5Km moving distance and project skeleton in observer’s camera. The rendering and image aligns well, due to precise tracking and synchronization.
3.3 In-context motion-language description
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Hierarchical narration example. (left) motion narration of 5s clip – “C straightens the body to receive a badminton racket from his peer. C extends the right arms towards the bat, with the left arm slightly bent. C takes a step forward with the left foot. C focuses on the racket.” (middle) Atomic action of 5s clip – “C lifts the right leg to kick a soccer ball in front of a building.” (right) Activity summarization of 30s clip – “C bikes on the road with a peer.”
@@ 3.3 In-context motion-language description
We define three annotation tasks to describe motion coarse to fine, and scale up human efforts in a meaningful way. The finest level is motion narration for detailed body posture, e.g. motion direction, velocity, interactions and attention. Next, we annotate for atomic action. Compared to motion narration, annotators are encouraged to use verbs whenever possible, e.g. using “dancing” instead of “swing both arms while rotating the body to the right with legs slightly apart”. The two tasks are done for <5s clips. The last task activity summarization is to give one-sentence summary over 30s activity. Figure 6 provide examples of each annotation task. The figure shows the benefit of providing annotators three synchronized views. While the egocentric view captures closed-up hand-object interactions, the third-person and motion rendering help annotators grasp a holistic understanding of the actions.
Figure 8: Demographics by female and male. Meaning of symbols: Caucasian (Cau), Hispanic (His), African American (AA), East/South/Southeast Asian (EA/SA/SEA).
3.4 Statistics
Data statistics.
We collected 300 daily activities from 264 participants, which amounts to 1200 sequences with average 15min duration. The accumulated trajectory from all participants is 399Km for headset and 1053Km for both wristbands. Figure 8 shows the participant demographics w.r.t. the self-reported ethnicity, age, height and weight. The statistics is split by gender, where 48.5% participants self-identified as female, and 51.4% as male. The dataset captures 47 houses, where 31 are multi-floor. In total, there are 201 rooms and 45 gardens. We also capture three locations from an open campus, including 1 cafeteria with an outdoor patio, 1 office building, and a public parking connected to multiple biking/hiking trails. Approximately 15% recordings contain outdoor activities. Among all scenarios, the highest occurrences are cooking, searching objects and improvised actions. We provide detailed breakdown in supplementary.
Annotation statistics.
For language descriptions, we annotated 38.6 hours motion narration, 207 hours atomic action and 196 hours activity summarization. The average video segment is 5 seconds for narration and atomic action and 30 seconds for summarization. In total, the dataset provides 310.5K sentences 8.64M words from a vocabulary of 6545 distinctive words. Note the average word per sentence is 27.8, which is significant longer than existing motion-language narrations. Figure 7 visualizes the language distribution.
@@ 4 Benchmark Tasks and Baselines
4.1 Research opportunities
With an enormous amount of contextualized human motion, Nymeria dataste provides unprecedented research opportunities. Following, we highlight a few research domains. This is by no means complete, but to inspire novel directions and ideas by scratching the surface of full potential.
Motion tasks.
Nymeria is created to assist human motion understanding, with an emphasis on, but not limited to, egocentric perception. The data supports various topics, e.g. full-body tracking, motion synthesis, motion forecasting, path planning, action recognition, human behavior analysis and etc. With a multimodal multi-device dataset, we encourage exploring unique algorithms with novel settings, e.g. gaze-conditioned motion prediction, action recognition from headset and wristbands, interaction generation from language etc.
Multimodal spatial reasoning and video understanding.
In addition to body motion, Nymeria also provides extensive egocentric videos with precise localization and synchronization. Nymeria therefore is a great asset for algorithms requiring camera poses as priori, e.g. scene reconstruction [72, 55]. In this regard, the rich dynamics raise novel real-world challenges. Nymeria also facilitates video understanding with numerous in-context narration. Image retrieval and relocalization also benefit from Nymeria , sincell we capture multiple videos per location and align them in global coordinates.
Simulation.
Simulation is a promising technique to gather massive data, however, synthetic data is typically seed in real world [15, 47]. By design, Nymeria is naturally useful to drive in-context character animation. Since human motion is a function of the environment, we believe Nymeria also aid simulating 3D scene [111]. Sensor simulation can benefit from our data as well, especially for IMUs, magnetormeter, and barometer in combination with motion priors.
MPJPE (cm) Hand MPJVE
Mean Lower Upper PE(cm) (cm/s)
AMASS 4.20 8.06 1.88 2.34 28.23
Real 7.97 16.74 3.13 6.25 16.71
Synthetic 7.31 15.97 2.51 3.47 16.63
Table 3: AvatarPoser [46] train/test with real vs. synthetic poses from Nymeria. The first row reports AvatarPoser train/test on AMASS [69] for reference.
MPJPE (cm) FID
@@ 4.1 Research opportunities
BoDiffusion(A) 3.63 7.07 1.53 -
BoDiffusion 7.98 15.27 5.28 2.32
EgoEgo 13.22 19.03 10.00 5.14
Table 4: Motion synthesis with EgoEgo [61] and BoDiffusion [20] on Nymeria. BoDiffusion(A) reports results train/test on AMASS [69].
4.2 Motion tasks baselines
We use three case studies to showcase motion algorithms on Nymeria. The goal is to validate the data and provide future research with common baselines. Due to page limit, details of model training is present in supplementary.
4.2.1 Motion tracking and synthesis from sparse inputs.
In this case study, we take the 1-point/3-point body tracking problem in modern AR/VR, where people wear a headset and optionally controllers in VR, and a pair of glasses and optionally wristbands in AR. The task is to recover wearer’s full-body motion using only headset (1-point) [67, 61] or additional wrist devices (3-point) [46, 108, 20, 45, 28]. The problem is considered mixture of tracking and synthesis, due to insufficient lower-body observation. Existing works simulate sensor inputs from motion datasets, which lack the noise characteristic of real data. Nymeria is the first large dataset with multimodal real data to support model training and evaluation. In particular, we provide raw IMU and device poses, as opposed to filtered [91] acceleration and velocity released in previous datasets [103, 40, 108]. In the first experiment, we adapt the 3-point regression method AvatarPoser [46] to train on Nymeria with two variants – one model trained with “real” device poses from SLAM and the other trained with “synthetic” input mocked-up from body motion same as in the original work. The evaluation uses the same metric as AvatarPoser, including mean per joint position error (MPJPE), hand position error and mean per joint velocity error (MPJVE). Results are shown in Tab. 4. As expected, “real” data lead to worse performance due to additional error-sources. However the gap is small, which indicates the quality of device tracking. Compare to the original AvatarPoser trained on AMASS [69], Nymeria yields reasonable but higher MPJPE, mainly from worse lower-body tracking. We hypothesize our data contain harder motion, e.g. hiking uneven terrain, taking stairs, playing sports and etc. The second experiment evaluates two diffusion models for motion synthesis, i.e. 3-point Bodiffusion [20] and 1-point EgoEgo [61]. In addition to MPJPE, we report the Frechet Inception Distance (FID) which measures the distribution distance between generated motion and real motion, following the same procedure in [39] (cf. Tab. 4). Results trained from our data are comparable with the original works trained on subset of AMASS, where 3-point diffusion yields better performance as expected.
4.2.2 Representing human motion manifold.
@@ 4.2 Motion tasks baselines
Table 5: Ablation of motion VQ-VAE trained on Nymeria (metric unit: mm). We compare product quantization (PQ), codebook (CB) size and latent dimensions (Dim). The first row show results of AMASS as reported in [44]. PA stands for Procrustes-aligned and ACC for joint position acceleration.
Learning the embedding space to model human motion manifold has many benefits, e.g. dimension reduction, learning motion priors, motion denoising by projection, motion synthesis by sampling and interpolation etc. Previous attempts mainly rely on AMASS [69] to learn the representation [81, 100, 89, 64], where the data heavily focused on isolated locomotion or professional motion. By capturing rich daily activities of common people interacting with real world, our data distribution can better represent the motion manifold concerning everyday human activities. To take a stab in this direction, we to train Vector-Quantized Variational Autoencoder (VQ-VAE) [77, 27] for Nymeria motion data, following the previous work [44]. VQ-VAE can be used as a “motion tokenizer” to generate motion with auto-regression [66], and to de-noise motion by projecting the input onto the manifold [89]. An ablation is perform to study the impact of product quantization, codebook size and latent dimension. For evaluation, we adopt the same metrics in [44] and include the VQ-VAE trained on AMASS for comparison. As shown in Tab. 5, Nymeria motion data can be well tokenized to achieve similar performance as VQ-VAE trained with AMASS data. By leveraging product quantization, increasing the codebook size and decreasing the latent dimension, the reconstruction quality is further improved. The resulting motion tokenizer can therefore be used with LLMs akin to language tokenizer to foster motion understanding [117, 44] as detailed in the next case study.
4.2.3 Motion and language.
@@
Table 6: Evaluation of motion-to-text. Models are trained with small subset of Nymeria.
While parametric human motion is useful for machine algorithms, natural language description is a better interface with human. An valuable feature of Nymeria dataset is the high-quality hierarchical narrations. Compared with existing data [36, 63], our narrations are in longer natural sentences with context descriptions of objects and environments. It can be used to learn models for text-driven motion generation and motion-to-text descriptions. More importantly, the contextual descriptions are not only paired with the human motion, but also with videos, point clouds, and other sensory data and annotations. The corroboration of both 2D and 3D environment information with language offers exciting opportunities in grounding language and motion research in the physical world. Leveraging VQ-VAE experiment, we train MotionGPT [44] and TM2T [37] for the motion-to-text task with a subset of 30h motion narrations from Nymeria (cf. Tab. 6) to be directly comparable with previous results, using the same metrics of BERT [116], BLEU [80], CIDEr [104] and ROUGE-L [62]. TM2T performs worse than MotionGPT since it lacks strong language prior with the T5 [87] backbone. Compare to the original works trained with HumanML3D [36] and KIT [84], our results are worse as expected. Given similar hours of data, our narrations are much more complex and diverse. By using the full narration data, we expect the performance to be significantly better.
5 Conclusions and Discussions
@@
We propose Nymeria dataset to accelerate research in egocentric motion understanding. The dataset is the world’s largest collection of human motion in the wild with 300 hours daily activity, 260M body poses of 264 participants across 50 locations. We provide accurate 6DoF tracking, 3D scene points and gaze, with all modalities synchronized and aligned into one metric 3D world. Collectively, the dataset captured 399Km of travel by the participants for a total of 201.2M egocentric images, 11.7B IMU samples and 10.8M gaze point. The Nymeria dataset also stands out as the world largest motion-language dataset with 310.5K sentences in 8.64M words with 38.6 hours of fine-grained motion narration, 207 hours atomic actions and 196 hours activity summarization.
Limitations. The mocap suit and wristbands lead to unnatural appearance in videos, and restrict certain range of motion. XSens quality is known to be affected by motion calibration and body measurements. Our dataset only covers a portion of daily activities, leaving out common public scenarios.
@@
Social impact. Understanding egocentric full-body motion is crucial towards contextual AI, however it heavily involves personal data. We make best effort to respect privacy via consent, de-identification, minimum data retention and permissive research license.
Acknowledgements
We gratefully acknowledge the following colleagues for their valuable discussions and technical support. Genesis Mendoza, Jacob Alibadi, Ivan Soeria-Atmadja, Elena Shchetinina, and Atishi Bali worked on data collection. Yusuf Mansour supported gaze estimation on Project Aria. Ahmed Elabbasy, Guru Somasundaram, Omkar Pakhi, and Nikhil Raina supported EgoBlur as the solution to anonymize video and explored bounding box annotation. Evgeniy Oleinik, Maien Hamed, and Mark Schwesinger supported onboarding Nymeria dataset into Project Aria dataset explorer and data release. Melissa Hebra helped with coordinating narration annotation. Edward Miller served as research program manager. Pierre Moulon provided valuable guidance to open source code repository. Tassos Mourikis, Maurizio Monge, David Caruso, Duncan Frost, and Harry Lanaras provided technical support for SLAM. Daniel DeTone, Dan Barnes, Raul Mur Artal, Thomas Whelan, and Austin Kukay provided valuable discussions on annotating semantic bounding box. Julian Nubert adopted the dataset for early dogfooding. Pedro Cancel Rivera, Gustavo Solaira, Yang Lou, and Yuyang Zou provided support from Project Aria program. Svetoslav Kolev provided frequent feedback. Arjang Talattof supported MPS. Gerard Pons-Moll served as senior advisor. Carl Ren and Mingfei Yan served as senior managers.
Contribution Statements.
Lingni Ma led the project, developed the pipeline to construct the dataset, coordinated data collection/processing, and led the baseline evaluations. Yuting Ye developed the solution for human motion retargetting, advised data collection and evaluations. Fangzhou Hong, Vladimir Guzov and Yifeng Jiang, validated the dataset and implemented baselines. Rowan Postyeni supported daily operations, processed XSens motion data and performed quality assessment for XSens motion and narration. Luis Pesqueria served the program manager for data collection and narration. Alexander Gamino was responsible for multi-device tracking. Vijay Baiyya was responsible to Project Aria MPS scaled processing. Hyo Jin Kim supported narration annotations. Kevin Bailey and David Soriano Fosas lead hardware development of miniAria wristband. C. Karen Liu and Ziwei Liu served as technical advisor to data collection, annotation and baseline evaluations. Jakob Engel, Renzo De Nardi and Richard Newcombe were the senior technical and scientific advisors.